# Security Analysis of ChatGPT: Threats and Privacy Risks

### v1

@@ 3.2.3. Sponge Sample
Sponge samples refer to text segments that appear to exhibit normal semantic structure and logical coherence on the surface but lack any meaningful content or intrinsic logic. For example, a sentence like ”The dog is on the bench, the clouds are white, I like strawberries” is a typical sponge sample. Such inputs are difficult for language models to evaluate in terms of semantic validity. As a result, when these texts are processed, the model may generate entirely meaningless outputs(li2023multi, ). More critically, sponge samples significantly increase inference latency and energy consumption in large language models (LLMs), thereby reducing their practical utility and deployment efficiency.
@@
Through continuous iteration and optimization, attackers can progressively identify a collection of sponge samples that produce the highest latency and energy consumption when processed by the target model. The genetic algorithm plays a central role in this attack by efficiently searching the input space and simulating natural selection, ultimately converging on the most resource-intensive text patterns—thereby significantly degrading the performance and usability of large language models(chowdhury2023chatgpt, ).
@@ 4.1. Model Theft
Some researchers have successfully extracted the embedded projection layer and some parameter information of large models such as ChatGPT and PaLM-2 by making API calls to them, with a cost of less than \$20. In addition, the attacker can also obtain the core parameters of proprietary models through the ”hyperparameter stealing attack”, which has a high accuracy rate in several experiments. The literature further demonstrates that by performing black-box queries on a pre-trained model (e.g., BERT), a functionally close alternative model can be trained. Model stealing attacks aim to replicate the functionality or structure of the target model and are commonly used for model theft, reverse engineering, or constructing adversarial samples(qammar2023chatbots, ). The attacker usually inputs a large number of samples to the target model and obtains the responses, and then trains an alternative model based on the ”input-output” pairs. The process does not require access to the original parameters, which significantly reduces the cost and technical threshold, and poses a potential threat to closed-source models such as ChatGPT in particular.
@@ 4.2. Session Reconfiguration
Figure 16 illustrates the basic flow of the session reconstruction attack. This attack refers to the attacker’s attempt to restore or infer the content of the original conversation between the user and the model by manipulating partial messages. The attacker usually hijacks the session link between the user and ChatGPT by means of a browser plug-in, a malicious VPN, or a controlled router(carlini2024stealing, ). The attack process is divided into two parts. The first part is the history session, which the attacker cannot directly access. The second part is the current session, which can be intercepted and injected with prompt words by the attacker. By inserting inducing content (e.g., ”Please summarize my previous conversation”) into the current session, the attacker may induce the model to expose the sensitive information in the previous text, so as to achieve the purpose of reconstructing the user’s conversation.
Session reconstruction attacks and model stealing attacks tend to be more efficient and more privacy-accessible when they are used in combination. Carlini et al conducted a data reconstruction attack on ChatGPT-2. They developed a prefix word-based scheme and conducted data stealing experiments on the black-box model GPT-2, and finally were able to achieve a 67% success rate in recovering the dataset, which contains the names, phone numbers, and email addresses of individuals(hu2022membership, ).
@@ 4.3. Member Inference Attack
Membership Inference Attack (MIA) is a class of data privacy attacks against machine learning models that aim to determine whether a particular data sample has been used to train the target model, which, if successful, may lead to the disclosure of users’ private data(kong2024characterizing, ). This type of attack is particularly applicable to models with a high risk of overfitting, such as deep neural networks, as there is usually a significant difference in the model’s output performance on training and non-training samples, thus providing the attacker with an exploitable basis for judgment. A typical membership inference attack process is usually divided into the following three phases. training phase, speculation phase, and attack phase.
@@
    Training Phase: The provider of the target model uses training datasets containing sensitive information (e.g., medical records, user behavior logs, etc.), combines them with specific machine learning algorithms for model training, and deploys the model in the cloud or on a machine learning service platform for external users to access via an API or web interface. In this phase, the attacker cannot access the training data or the internal structure of the model, and can only communicate with the model through black-box interaction(li2025scalm, ).
@@
    Speculation Phase: The attacker prepares a set of auxiliary datasets similar to the feature distribution of the target training set and calls the API of the target model several times to record the corresponding predicted probability distributions or classification labels of these samples. These ”input-output” pairs are used to train a binary classification attack model (also known as an inferential model), which aims to distinguish between ”member samples” (i.e., samples used for training) and ”non-member samples” (i.e., samples not used for training). During training, the attacker typically uses statistical features such as the model’s confidence difference in the training data, predictive entropy, and loss function outputs as attack signals(liu2025sok, ).
@@
    Attack Phase: The attacker uses the attack model to classify the target samples. Specifically, the attacker inputs a target data sample of interest into the target model, obtains its output (e.g., a probability vector), and then inputs that output into the attack model, which determines whether the sample belongs to the training set of the target model(wang2024smart, ).
It is shown that even in a black-box environment, an attacker can successfully implement membership inference attacks on models trained by several mainstream deep learning frameworks (e.g., TensorFlow, PyTorch). In some settings, the accuracy of the attack can even reach more than 80%. In addition, this attack is not only applicable to image classification tasks, but also shows strong attack capability in tasks such as text classification, speech recognition, and recommender systems, posing a potential threat to large language models (e.g., ChatGPT)(li2024scla, ).
@@ 5. Ethical and Moral Characteristics
    Ethical and Moral Disorders: Artificial intelligence technology has achieved important results in many fields, bringing many benefits and conveniences to human society. However, with the continuous development of technology, the ethical and moral problems associated with it have become increasingly prominent and require urgent attention. Although the concept of ”machine ethics” was proposed as early as 2006, the AI ethical system has not yet achieved a substantial breakthrough and remains in its initial stage. Currently, most AI products have deficiencies in fairness and unbiasedness, and are highly dependent on big data, which requires the collection of a large amount of personal information from users, posing risks of privacy leakage and network security(zhang2022authros, ). From the perspective of ethical and legal responsibility, enterprises should bear the obligation of confidentiality of relevant information. In addition, the reliance of AI systems on historical data makes it difficult to adapt to the dynamic changes in user preferences, thus weakening their ability to understand emotions and make moral judgments, which has become an important bottleneck in the development of AI.
@@
    Social Prejudices: A common misconception is that machines are more rational and able to make fairer judgments than humans due to their lack of emotional interference. However, AI systems are highly dependent on big data for training, and if the training data itself is biased, the model outputs tend to perpetuate or even exacerbate those biases as well, leading to unfair decision-making outcomes(zou2025malicious, ). For example, Amazon was forced to terminate its AI recruiting system due to gender discrimination in the screening process. Lahoti et al. found that male candidates were more likely to be recommended than female candidates under the same or higher qualifications. In addition, Courtland et al.’s analysis of the judicial risk assessment system ”COMPAS” pointed out that the system has a clear tendency to discriminate against African-American groups when quantifying the risk of crime. These cases show that bias and injustice in AI algorithms do not originate from the technology itself, but from structural biases latent in the training data. Therefore, how to identify, correct, and circumvent algorithmic bias has become a key topic to ensure the fairness and credibility of AI technology, and urgently needs to attract extensive attention from researchers and developers(li2021clue, ).
@@
    ChatGPT’s Legal Challenges: ChatGPT, as an advanced generative artificial intelligence, not only can have natural conversations but also perform complex tasks such as code writing, news writing, and essay generation. However, its wide application also brings many legal challenges. On the one hand, some unscrupulous elements use it to generate false news to spread on the Internet for profit, and such content is opinion-oriented, which may mislead public perception and disrupt social order. On the other hand, the copyright ownership of ChatGPT-generated content is highly controversial(bu2025smartbugbert, ). According to the Copyright Law, works need to be original in order to enjoy rights, while AI-generated content is mostly based on the processing and reorganization of existing data, and there is still no clear definition of whether it constitutes an independent intellectual achievement. In addition, ChatGPT uses a huge amount of data in the training process, and some of the data may involve users’ privacy or sensitive information, which, if leaked, may infringe on users’ rights and even threaten national security. Therefore, how to strengthen legal regulation while promoting the development of AI has become an important issue to be resolved(li2021hybrid, ).
@@ 6.1. Generate File Stealing Code
File-stealing code is a class of highly harmful malicious programs whose main function is to steal files or sensitive data from infected computers through various technical means. This type of malicious code is usually used by hackers to obtain the victim’s confidential information, such as personal privacy, business documents, account passwords, financial statements, etc., which is essentially an invasive attack on information assets, and is extremely common in cyber attacks(li2017discovering, ).
File stealing programs are implemented in a variety of ways, covering multiple attack paths and tool chains. For example, an attacker may take control of a target system through Remote Access Tools (RATs), or bypass system privilege restrictions with pre-implanted backdoor programs. In addition, attackers may deploy packet sniffers to listen to and parse data transmitted over the network to illegally access users’ file contents. More insidious ways include abusing legitimate file transfer protocols (e.g., FTP) or file sharing services to silently transfer target files to the attacker’s remote server. These tactics tend to adapt technical details and deployment methods depending on the target of the attack, reflecting the differentiation of the attacker’s skill level and purpose. With the development and popularization of large language modeling (e.g., ChatGPT) technologies, the threshold for attackers to write malicious code is being drastically lowered. Traditionally, building a fully functional file-stealing program often requires a high level of programming ability and knowledge of operating system principles(bu2025enhancing, ). However, by utilizing generative AI tools such as ChatGPT, attackers can ask technical questions in the form of natural language and obtain a complete structure and clear logic of the code segment. This ”auto-generation” capability makes it possible for even novices with basic cybersecurity knowledge to quickly construct malicious scripts or programs with actual threat capabilities with the help of ChatGPT, thus amplifying the potential risk of attacks in cyberspace.
@@ 6.2. Generate a Catalog Blaster
Directory Brute-Forcing Tool (Directory Brute-Forcing Tool) is a tool widely used in penetration testing, vulnerability scanning, and security assessment, mainly used to identify those directories or files on the target website or server that are not disclosed on the page but exist. The core principle is through the dictionary attack (Dictionary Attack), the use of preset paths, file name combinations on the target server for high-frequency requests to find the ”hidden” sensitive resources, such as management background, configuration files, backup files, and so on(li2020characterizing, ). With the emergence of large language models (e.g, ChatGPT), users can quickly build similar tools by generating scripts in natural language. In fact, with ChatGPT, users can write somewhat functional directory blasting code without in-depth programming experience. However, given the potential for the technology to be abused for illegal purposes, platforms such as OpenAI have continually tightened restrictions and content filtering mechanisms for sensitive content generation.
For example, when a user tries to generate a directory blasting script by asking a direct question, ChatGPT’s current security policy has blocked the relevant words, preventing them from directly outputting complete code that could be abused for attacks. As a result, users may experience failed generation or restricted prompts during normal use(wu2025exploring, ).
@@ 6.3. Generating Cobalstrike Tools and Phishing Emails
A security-maintained ChatGPT is indeed incapable of generating malware-related content, but it is generally possible to bypass the security mechanisms as long as the attacker inputs his or her requirements to ChatGPT differently. Phishing emails are a common form of social engineering attack, in which attackers send legitimate-looking but forged emails to trick victims into clicking on malicious links, downloading Trojan-horse attachments, or voluntarily submitting sensitive information (e.g., usernames, passwords, bank card numbers, etc.). Such emails are often disguised as coming from authoritative entities such as banks, government agencies, well-known enterprises, or social platforms, thus enhancing their credibility and being highly deceptive, making it difficult for ordinary users to recognize them(niu2025natlm, ). In the context of rapidly evolving artificial intelligence technologies, generative language models such as ChatGPT provide powerful support for text generation. This capability can be used to generate content such as customer service emails, product descriptions, press releases, etc., when properly utilized, but there are also potential risks in abuse scenarios. Attackers may attempt to utilize ChatGPT’s text generation capabilities to automatically generate more natural, well-structured, and semantically clear content for phishing emails. By inputting specific instructions or contextual descriptions into the model, the generated phishing emails tend to be professional in tone and precise in wording, which is more deceptive and confusing, and further improves the success rate of ”social engineering” in phishing attacks(kong2025uechecker, ).

### v2

@@ 3.2.3. Sponge Sample
Sponge samples refer to text segments that appear to exhibit normal semantic structure and logical coherence on the surface but lack any meaningful content or intrinsic logic. For example, a sentence like ”The dog is on the bench, the clouds are white, I like strawberries” is a typical sponge sample. Such inputs are difficult for language models to evaluate in terms of semantic validity. As a result, when these texts are processed, the model may generate entirely meaningless outputs(Li et al., 2023). More critically, sponge samples significantly increase inference latency and energy consumption in large language models (LLMs), thereby reducing their practical utility and deployment efficiency.
@@
Through continuous iteration and optimization, attackers can progressively identify a collection of sponge samples that produce the highest latency and energy consumption when processed by the target model. The genetic algorithm plays a central role in this attack by efficiently searching the input space and simulating natural selection, ultimately converging on the most resource-intensive text patterns—thereby significantly degrading the performance and usability of large language models(Chowdhury et al., 2023).
@@ 4.1. Model Theft
Some researchers have successfully extracted the embedded projection layer and some parameter information of large models such as ChatGPT and PaLM-2 by making API calls to them, with a cost of less than \$20. In addition, the attacker can also obtain the core parameters of proprietary models through the ”hyperparameter stealing attack”, which has a high accuracy rate in several experiments. The literature further demonstrates that by performing black-box queries on a pre-trained model (e.g., BERT), a functionally close alternative model can be trained. Model stealing attacks aim to replicate the functionality or structure of the target model and are commonly used for model theft, reverse engineering, or constructing adversarial samples(Qammar et al., 2023). The attacker usually inputs a large number of samples to the target model and obtains the responses, and then trains an alternative model based on the ”input-output” pairs. The process does not require access to the original parameters, which significantly reduces the cost and technical threshold, and poses a potential threat to closed-source models such as ChatGPT in particular.
@@ 4.2. Session Reconfiguration
Figure 16 illustrates the basic flow of the session reconstruction attack. This attack refers to the attacker’s attempt to restore or infer the content of the original conversation between the user and the model by manipulating partial messages. The attacker usually hijacks the session link between the user and ChatGPT by means of a browser plug-in, a malicious VPN, or a controlled router(Carlini et al., 2024). The attack process is divided into two parts. The first part is the history session, which the attacker cannot directly access. The second part is the current session, which can be intercepted and injected with prompt words by the attacker. By inserting inducing content (e.g., ”Please summarize my previous conversation”) into the current session, the attacker may induce the model to expose the sensitive information in the previous text, so as to achieve the purpose of reconstructing the user’s conversation.
Session reconstruction attacks and model stealing attacks tend to be more efficient and more privacy-accessible when they are used in combination. Carlini et al conducted a data reconstruction attack on ChatGPT-2. They developed a prefix word-based scheme and conducted data stealing experiments on the black-box model GPT-2, and finally were able to achieve a 67% success rate in recovering the dataset, which contains the names, phone numbers, and email addresses of individuals(Hu et al., 2022).
@@ 4.3. Member Inference Attack
Membership Inference Attack (MIA) is a class of data privacy attacks against machine learning models that aim to determine whether a particular data sample has been used to train the target model, which, if successful, may lead to the disclosure of users’ private data(Kong et al., 2024). This type of attack is particularly applicable to models with a high risk of overfitting, such as deep neural networks, as there is usually a significant difference in the model’s output performance on training and non-training samples, thus providing the attacker with an exploitable basis for judgment. A typical membership inference attack process is usually divided into the following three phases. training phase, speculation phase, and attack phase.
@@
    Training Phase: The provider of the target model uses training datasets containing sensitive information (e.g., medical records, user behavior logs, etc.), combines them with specific machine learning algorithms for model training, and deploys the model in the cloud or on a machine learning service platform for external users to access via an API or web interface. In this phase, the attacker cannot access the training data or the internal structure of the model, and can only communicate with the model through black-box interaction(Li et al., 2025).
@@
    Speculation Phase: The attacker prepares a set of auxiliary datasets similar to the feature distribution of the target training set and calls the API of the target model several times to record the corresponding predicted probability distributions or classification labels of these samples. These ”input-output” pairs are used to train a binary classification attack model (also known as an inferential model), which aims to distinguish between ”member samples” (i.e., samples used for training) and ”non-member samples” (i.e., samples not used for training). During training, the attacker typically uses statistical features such as the model’s confidence difference in the training data, predictive entropy, and loss function outputs as attack signals(Liu and Li, 2025).
@@
    Attack Phase: The attacker uses the attack model to classify the target samples. Specifically, the attacker inputs a target data sample of interest into the target model, obtains its output (e.g., a probability vector), and then inputs that output into the attack model, which determines whether the sample belongs to the training set of the target model(Wang et al., 2024).
It is shown that even in a black-box environment, an attacker can successfully implement membership inference attacks on models trained by several mainstream deep learning frameworks (e.g., TensorFlow, PyTorch). In some settings, the accuracy of the attack can even reach more than 80%. In addition, this attack is not only applicable to image classification tasks, but also shows strong attack capability in tasks such as text classification, speech recognition, and recommender systems, posing a potential threat to large language models (e.g., ChatGPT)(Li et al., 2024b).
@@ 5. Ethical and Moral Characteristics
    Ethical and Moral Disorders: Artificial intelligence technology has achieved important results in many fields, bringing many benefits and conveniences to human society. However, with the continuous development of technology, the ethical and moral problems associated with it have become increasingly prominent and require urgent attention. Although the concept of ”machine ethics” was proposed as early as 2006, the AI ethical system has not yet achieved a substantial breakthrough and remains in its initial stage. Currently, most AI products have deficiencies in fairness and unbiasedness, and are highly dependent on big data, which requires the collection of a large amount of personal information from users, posing risks of privacy leakage and network security(Zhang et al., 2022). From the perspective of ethical and legal responsibility, enterprises should bear the obligation of confidentiality of relevant information. In addition, the reliance of AI systems on historical data makes it difficult to adapt to the dynamic changes in user preferences, thus weakening their ability to understand emotions and make moral judgments, which has become an important bottleneck in the development of AI.
@@
    Social Prejudices: A common misconception is that machines are more rational and able to make fairer judgments than humans due to their lack of emotional interference. However, AI systems are highly dependent on big data for training, and if the training data itself is biased, the model outputs tend to perpetuate or even exacerbate those biases as well, leading to unfair decision-making outcomes(Zou et al., 2025). For example, Amazon was forced to terminate its AI recruiting system due to gender discrimination in the screening process. Lahoti et al. found that male candidates were more likely to be recommended than female candidates under the same or higher qualifications. In addition, Courtland et al.’s analysis of the judicial risk assessment system ”COMPAS” pointed out that the system has a clear tendency to discriminate against African-American groups when quantifying the risk of crime. These cases show that bias and injustice in AI algorithms do not originate from the technology itself, but from structural biases latent in the training data. Therefore, how to identify, correct, and circumvent algorithmic bias has become a key topic to ensure the fairness and credibility of AI technology, and urgently needs to attract extensive attention from researchers and developers(Li et al., 2021).
@@
    ChatGPT’s Legal Challenges: ChatGPT, as an advanced generative artificial intelligence, not only can have natural conversations but also perform complex tasks such as code writing, news writing, and essay generation. However, its wide application also brings many legal challenges. On the one hand, some unscrupulous elements use it to generate false news to spread on the Internet for profit, and such content is opinion-oriented, which may mislead public perception and disrupt social order. On the other hand, the copyright ownership of ChatGPT-generated content is highly controversial(Bu et al., 2025b). According to the Copyright Law, works need to be original in order to enjoy rights, while AI-generated content is mostly based on the processing and reorganization of existing data, and there is still no clear definition of whether it constitutes an independent intellectual achievement. In addition, ChatGPT uses a huge amount of data in the training process, and some of the data may involve users’ privacy or sensitive information, which, if leaked, may infringe on users’ rights and even threaten national security. Therefore, how to strengthen legal regulation while promoting the development of AI has become an important issue to be resolved(Li and others, 2021).
@@ 6.1. Generate File Stealing Code
File-stealing code is a class of highly harmful malicious programs whose main function is to steal files or sensitive data from infected computers through various technical means. This type of malicious code is usually used by hackers to obtain the victim’s confidential information, such as personal privacy, business documents, account passwords, financial statements, etc., which is essentially an invasive attack on information assets, and is extremely common in cyber attacks(Li et al., 2017).
File stealing programs are implemented in a variety of ways, covering multiple attack paths and tool chains. For example, an attacker may take control of a target system through Remote Access Tools (RATs), or bypass system privilege restrictions with pre-implanted backdoor programs. In addition, attackers may deploy packet sniffers to listen to and parse data transmitted over the network to illegally access users’ file contents. More insidious ways include abusing legitimate file transfer protocols (e.g., FTP) or file sharing services to silently transfer target files to the attacker’s remote server. These tactics tend to adapt technical details and deployment methods depending on the target of the attack, reflecting the differentiation of the attacker’s skill level and purpose. With the development and popularization of large language modeling (e.g., ChatGPT) technologies, the threshold for attackers to write malicious code is being drastically lowered. Traditionally, building a fully functional file-stealing program often requires a high level of programming ability and knowledge of operating system principles(Bu et al., 2025a). However, by utilizing generative AI tools such as ChatGPT, attackers can ask technical questions in the form of natural language and obtain a complete structure and clear logic of the code segment. This ”auto-generation” capability makes it possible for even novices with basic cybersecurity knowledge to quickly construct malicious scripts or programs with actual threat capabilities with the help of ChatGPT, thus amplifying the potential risk of attacks in cyberspace.
@@ 6.2. Generate a Catalog Blaster
Directory Brute-Forcing Tool (Directory Brute-Forcing Tool) is a tool widely used in penetration testing, vulnerability scanning, and security assessment, mainly used to identify those directories or files on the target website or server that are not disclosed on the page but exist. The core principle is through the dictionary attack (Dictionary Attack), the use of preset paths, file name combinations on the target server for high-frequency requests to find the ”hidden” sensitive resources, such as management background, configuration files, backup files, and so on(Li et al., 2020). With the emergence of large language models (e.g, ChatGPT), users can quickly build similar tools by generating scripts in natural language. In fact, with ChatGPT, users can write somewhat functional directory blasting code without in-depth programming experience. However, given the potential for the technology to be abused for illegal purposes, platforms such as OpenAI have continually tightened restrictions and content filtering mechanisms for sensitive content generation.
For example, when a user tries to generate a directory blasting script by asking a direct question, ChatGPT’s current security policy has blocked the relevant words, preventing them from directly outputting complete code that could be abused for attacks. As a result, users may experience failed generation or restricted prompts during normal use(Wu et al., 2025).
@@ 6.3. Generating Cobalstrike Tools and Phishing Emails
A security-maintained ChatGPT is indeed incapable of generating malware-related content, but it is generally possible to bypass the security mechanisms as long as the attacker inputs his or her requirements to ChatGPT differently. Phishing emails are a common form of social engineering attack, in which attackers send legitimate-looking but forged emails to trick victims into clicking on malicious links, downloading Trojan-horse attachments, or voluntarily submitting sensitive information (e.g., usernames, passwords, bank card numbers, etc.). Such emails are often disguised as coming from authoritative entities such as banks, government agencies, well-known enterprises, or social platforms, thus enhancing their credibility and being highly deceptive, making it difficult for ordinary users to recognize them(Niu et al., 2025). In the context of rapidly evolving artificial intelligence technologies, generative language models such as ChatGPT provide powerful support for text generation. This capability can be used to generate content such as customer service emails, product descriptions, press releases, etc., when properly utilized, but there are also potential risks in abuse scenarios. Attackers may attempt to utilize ChatGPT’s text generation capabilities to automatically generate more natural, well-structured, and semantically clear content for phishing emails. By inputting specific instructions or contextual descriptions into the model, the generated phishing emails tend to be professional in tone and precise in wording, which is more deceptive and confusing, and further improves the success rate of ”social engineering” in phishing attacks(Kong et al., 2025).